# Feasibility-Aware Decision-Focused Learning for Predicting Parameters in the Constraints

### v1

@@ Abstract
When some parameters of a constrained optimization problem (COP) are uncertain, this gives rise to a predict-then-optimize (PtO) problem, comprising two stages – the prediction of the unknown parameters from contextual information and the subsequent optimization using those predicted parameters. Decision-focused learning (DFL) implements the first stage by training a machine learning (ML) model to optimize the quality of the decisions made using the predicted parameters. When parameters in the constraints of a COP are predicted, the predicted parameters can lead to infeasible solutions. Therefore, it is important to simultaneously manage both feasibility and decision quality. We develop a DFL framework for predicting constraint parameters in a generic COP. While prior works typically assume that the underlying optimization problem is a linear program (LP) or integer linear program (ILP), our approach makes no such assumption. We derive two novel loss functions based on maximum likelihood estimation (MLE): the first one penalizes infeasibility (by penalizing when the predicted parameters lead to infeasible solutions), and the second one penalizes suboptimal decisions (by penalizing when the true optimal solution is infeasible under the predicted parameters). We introduce a single tunable parameter to form a weighted average of the two losses, allowing decision-makers to balance suboptimality and feasibility. We experimentally demonstrate that adjusting this parameter provides a decision-maker the control over the trade-off between the two. Moreover, across several COP instances, we find that for a single value of the tunable parameter, our method matches the performance of the existing baselines on suboptimality and feasibility.
@@ 1 Introduction
Many real-world optimization problems involve parameters that are unknown at decision time – for example, uncertain customer demand in manufacturing or unknown traffic in delivery routing. This gives rises to the predict-then-optimize (PtO) [18] problems, where a machine learning (ML) model first makes point predictions of the unknown parameters using contextual features, and then the optimization problem is solved using these predictions. In a prediction-focused learning (PFL) approach, the predictive model is trained solely to improve the accuracy of parameter predictions minimizing standard ML losses. By doing so, PFL does not consider the effect of prediction errors on the solution to the optimization problem. For example, consider a portfolio allocation problem under budget constraints, where the goal is to predict the future bid prices of stocks. If the bid price of a stock is higher than the available budget, the customer will not buy it. Hence, overestimating the bid price of that stock would change the portfolio allocation. However, underestimating it could change the allocation. Therefore, PFL does not account for how different types of prediction errors influence the quality of the final decision. To address this shortcoming of PFL, decision-focused learning (DFL) directly trains the ML model to minimize a task-specific decision loss that reflects the quality of the solution produced by using the predicted parameters. With a slight abuse of notation, we will use the term ‘predicted solution’ to denote the solution produced by using the predicted parameters, although the ML model does not directly predict the solution, it predicts parameters that are passed to an optimization solver at inference time.
@@
Most existing DFL works [10, 28, 19, 11] focus on cases where the predicted parameters appear only in the objective function. In such settings, decision loss is often measured by regret, which measures suboptimality. In contrast, we focus on predicting parameters that appear in the constraints of the optimization problem. In this setting too, suboptimal solutions may arise because the feasible space defined by the predicted parameters can exclude the true optimal solution. However, focusing only on suboptimality is insufficient. It can also be that the predicted solution is infeasible with respect to the feasible space of the true parameters. Therefore, the ML model should also be trained to minimize the likelihood of such outcomes.
@@
Focusing only on suboptimality or infeasibility can lead to unrealistic predictions that are not useful in practice. To address this, Hu et al. [12] model decision-making with uncertain constraint parameters as a two-stage process: first, a solution is generated using the predicted parameters; then, if needed, corrective actions – incurring additional cost – are carried out to improve infeasibility or suboptimality. Subsequent works [13, 14] have extended this line of research.
@@
One limitation of their approach is that the corrective action is often problem-specific, and modeling it as a specific form of optimization problem, as they have done, may not be realistic. To avoid the complexity of a second stage, we instead focus on predicting parameters that minimize both infeasbility and suboptimality in the first-stage solution itself. To this end, we first propose two novel loss functions – Infeasibility Penalty Loss (IPL) and Optimality-Preserving Loss (OPL). The motivation behind IPL is that if solving with the predicted parameters leads to a solution that is infeasible according to the true parameters, then such infeasibility should be (proportionally) penalized. On the other hand, the OPL loss will penalize if the predicted parameters make the true optimal solution infeasible. This would avoid obtaining a suboptimal solution when solving with the predicted parameters.
@@
These two losses often contradict each other, because IPL encourages strict or tight parameter predictions that ensure the predicted parameters lead to a truly feasible solution, while OPL incentivizes loose parameter predictions to ensure the optimal solution is not cut away. So, these two losses have opposing goals. Moreover, in practice, a decision-maker may favor more feasibility or optimality. To support this subjective preference of one loss over the other, we propose to minimize a weighted average of the two losses. In this way we introduce a tunable weighting mechanism that adjusts the total loss based on the decision-maker’s preference. Our experiments show that this approach results in decisions that better reflect preferences of the decision-maker.
@@
We compare our proposed technique with the PFL approach, which minimizes the mean squared error (MSE) loss, and with some existing DFL approaches. These approaches do not allow the decision-maker to adjust the predicted solution based on their subjective preference between optimality and feasibility. Moreover, the DFL approaches for constraint prediction are limited to linear programs (LPs) or integer or mixed-integer LPs (ILPs and MILPs), or they focus on predicting full constraint systems rather than individual parameters within the constraints. Our experimental results across various optimization problem instances show that, for a single value of the tunable parameter, our method matches or outperforms existing baselines in terms of both suboptimality and feasibility. More importantly, our approach offers a simple way for decision-makers to control the trade-off between suboptimality and infeasibility using a single parameter.
@@ 2 Related Literature
Differentiating through optimization solver.
@@
Differentiating through optimization problems has gained attention in ML, as it allows solvers to be embedded into gradient-based training. This is often done by ADMM [7, 25] or by implicit differentiation [6] of the optimality conditions. Implicit differentiation techniques have been developed for quadratic programs [2] and cone programs [1]. However, for combinatorial problems like ILPs, the true gradient is zero almost everywhere because small changes in the parameters usually do not affect the optimal solution. To address this, Paulus et al. [22] propose Comboptnet, a technique for ILPs to compute pseudo-gradients based on the Euclidean distance between the solution and the constraints.
Decision-focused learning.
DFL trains ML models to directly minimize decision loss, which requires backpropagating through the optimization problem. Most DFL works focus on predicting parameters in the objective, so they primarily focus on computing gradients of the objective parameters. Some works in this field are based on differentiating through the COP after smoothing it, either with a regularization [28, 16, 11] or a random perturbation [3, 24, 21, 8]. Another approach is to define informative pseudogradients of the decision loss [10, 17, 19]. We refer readers to a recent survey paper [18] for a more comprehensive overview.
Predicting constraints.
@@
Only a few DFL works focus on predicting constraint parameters. Most of these [12, 13, 14] model decision-making as a two-stage process, allowing corrective actions when predictions lead to infeasible or suboptimal solutions. Our approach instead treats this as a single-step decision problem with no corrective actions. Another approach [9] simultaneously learns the objective and the constraints, but the constraints are soft and the problem must be formulated as a discrete graphical model. The work by Nandwani et al. [20] is the closest to ours. However, their aim is to a learn the full set of constraints from the features, whereas we focus on learning only the parameters within a known constraint structure. Their method does not easily generalize to our setting, in which the learned parameters must remain compatible with the known constraints to make good quality solutions. Moreover, both Hu et al. [13] and Nandwani et al. [20] restrict their methods to LPs and ILPs, while our framework is more general and applicable to a broader class of optimization problems. The prediction of constraints is also central in constraint acquisition, for satisfiability [15] or constraint satisfaction [4, 27] problems, but the setting differs: the structure of the constraints is unknown and the prediction is non-contextual (it does not depend on any input features).
@@ 3 Preliminaries
We will consider a constrained optimization problem (COP) in the following form:
@@
where $\Omega$ defines the domain of the decision variable vector $\bm{x}$. $\Omega$ may be a subset of real numbers or integers, depending on the type of optimization problems and determine the upper and lower bounds of the decision variables. We will assume that the constraint functions $g_{i}(\bm{x};\bm{\rho})$ are differentiable functions. We do not explicitly express equality constraints, as a constraint in the form of $h_{j}(\bm{x};\bm{\rho})=0$, can be expressed by two inequality constraints $h_{j}(\bm{x};\bm{\rho})\leq 0$ and $h_{j}(\bm{x};\bm{\rho})\leq 0$. To make the notation simpler, we will also denote the set of constraints as $\mathcal{S}$. Further, we use $\mathcal{F}(\rho)$ to denote the feasible set when the constraint parameters take the value $\rho$. More formally,
@@
We use $\bm{x}^{\star}(\bm{q},\bm{\rho})$ to denote a (possibly non-unique) solution to the optimization problem for a given objective parameter $\bm{q}$ and constraint parameter $\bm{\rho}$. When multiple optimal solutions exist, we define the set of all optimal solutions as $\bm{W}(\bm{q},\bm{\rho})=\{\bm{x}:f(\bm{x};\bm{q})\leq f(\bm{x}^{\prime};\bm{q});\;\forall\bm{x}^{\prime}\in\mathcal{F}(\bm{\rho})\}$. Note that $\bm{W}(\bm{q},\bm{\rho})$ is a subset of the feasible region $\mathcal{F}(\bm{\rho})$. So, if $\mathcal{F}$ is empty (i.e., no feasible solution), $\bm{W}$ is also empty. In the special case of a unique optimal, $\bm{W}(\bm{q},\bm{\rho})$ is a singleton. Our implementation will assume the solution is unique, as we rely on a solver to solve the COP and they return one optimal solution. When multiple optimal solutions exist, the commercial solvers we use return a single optimal solution by breaking ties using a predefined rule.
@@
We consider the multi-dimensional 0-1 knapsack problem (MDKP)[23]. The objective of this problem is to choose a maximal value subset from a given set of items, subject to $M$ capacity constraints. Let $\rho_{i}$ be the capacity in dimension $i$. There are $\mathit{N}$ number of items, the value of each item is $q_{n}$, and $\rho_{(n,i)}$ is the weight of the item $n$ in dimension $i$. Each decision variable, $x_{n}$, can be either zero or one; hence $\Omega=\{0,1\}^{N}$. This optimization can be modeled as an ILP as follows:
Here, $\mathcal{S}$ consists of the $M$ capacity constraints, to be predicted.
@@
We will develop a framework based on maximum likelihood estimation to predict constraint parameters. Our goal is to maximize the log-likelihood of the observed data $\mathcal{D}$. In PFL, the parameter $\theta$ would be estimated to increase the likelihood of the observation, $\bm{\rho}_{\kappa}$, given the feature, $\bm{\phi}_{\kappa}$. However, in DFL training, the focus is on the likelihood of $\bm{x}^{\star}(\bm{q}_{\kappa},\bm{\rho}_{\kappa})$ being the optimal solution, given $\bm{q}_{\kappa}$ and $\bm{\phi}_{\kappa}$. To put in another way, DFL with an MLE perspective consists of estimating $\theta$ to predict the constraint parameter vector, $\bm{\hat{\rho}}_{\kappa}$, so that $\bm{x}^{\star}(\bm{q}_{\kappa},\bm{\rho}_{\kappa})$ is the most likely solution to the parameter, $\bm{\hat{\rho}}_{\kappa}=\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})$.
@@ 4 Methodology
We start by defining a function $\mathit{UNSAT}_{i}(\bm{x},\bm{\rho}_{\kappa})$ to indicate whether the $i$-th constraint is violated by an assignment, $\bm{x}$, when the constraint parameter is set to the vector $\bm{\rho}_{\kappa}$. It is defined as:
@@
As we want to maximize the likelihood of leading to a good solution, we have to model the probability of violating the constraints using the predicted parameters. Given a predicted parameter, $\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})$, we model the parametric probability, $\mathit{P}_{\theta}\big(\mathit{UNSAT}_{i}(\bm{x},\mathcal{M}_{\theta}(\bm{\phi}_{\kappa}))\big|\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\big)$, in the following manner:
@@
As, $\mathit{P}_{\theta}\left(\mathit{SAT}_{i}\left(\bm{x},\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\right)\middle|\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\right)$ = 11- $\mathit{P}_{\theta}\left(\mathit{UNSAT}_{i}\left(\bm{x},\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\right)\middle|\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\right)$, we can write:
@@
4.1 Loss Function to Penalize for Infeasibility-Inducing Predictions
We first propose a loss function that penalizes the predicted parameters when the resulting solution violates the true constraints. Let $\bm{x}_{\mathit{neg}}$ be an assignment that is infeasible under the true parameters. As it is easy to determine which constraints are violated by $\bm{x}_{\mathit{neg}}$, we develop an MLE framework to maximize the likelihood of violating those constraints under the predicted parameters. Given a $\bm{x}_{\mathit{neg}}$, the log-likelihood of violating these constraints can be written in the following form:

### v2

@@ Abstract
When some parameters of a constrained optimization problem (COP) are uncertain, this gives rise to a predict-then-optimize (PtO) problem, comprising two stages: the prediction of the unknown parameters from contextual information and the subsequent optimization using those predicted parameters. Decision-focused learning (DFL) implements the first stage by training a machine learning (ML) model to optimize the quality of the decisions made using the predicted parameters. When the predicted parameters occur in the constraints, they can lead to infeasible solutions. Therefore, it is important to simultaneously manage both feasibility and decision quality. We develop a DFL framework for predicting constraint parameters in a generic COP. While prior works typically assume that the underlying optimization problem is a linear program (LP) or integer LP (ILP), our approach makes no such assumption. We derive two novel loss functions based on maximum likelihood estimation (MLE): the first one penalizes infeasibility (by penalizing predicted parameters that lead to infeasible solutions), while the second one penalizes suboptimal decisions (by penalizing predicted parameters that make the true optimal solution infeasible). We introduce a single tunable parameter to form a weighted average of the two losses, allowing decision-makers to balance suboptimality and feasibility. We experimentally demonstrate that adjusting this parameter provides decision-makers control over this trade-off. Moreover, across several COP instances, we show that adjusting the tunable parameter allows a decision-maker to prioritize either suboptimality or feasibility, outperforming the performance of existing baselines in either objective.
@@ 1 Introduction
Many real-world optimization problems involve parameters that are unknown at decision time, such as uncertain customer demand in manufacturing problems or unknown traffic in delivery routing. This gives rise to predict-then-optimize (PtO) problems [22], where a machine learning (ML) model first makes point predictions of the unknown parameters from contextual features, after which the resulting instantiated optimization problem is solved.
In prediction-focused learning (PFL), the predictive model is trained solely to maximize the accuracy of parameter predictions by minimizing standard ML losses. By doing so, PFL does not consider the effect of prediction errors on the solution to the optimization problem. For example, consider a vehicle routing problem, where one first needs to predict unknown customer demands, to then make routing decisions. In this setting, overestimations and underestimations can have differing effects. Overestimating demand may lead to a conservative suboptimal decision, while underestimating demand may lead to an infeasible decision. However, these differing effects are not accounted for by PFL losses.
To address this shortcoming of PFL, decision-focused learning (DFL) directly trains the ML model to minimize a task-specific decision loss that reflects the quality of the solution produced by using the predicted parameters. With a slight abuse of notation, we will use the term ‘predicted solution’ to denote the solution produced by using the predicted parameters, although the ML model does not directly predict the solution, it predicts parameters that are passed to an optimization solver at inference time.
@@
Most existing DFL works [12, 13, 23, 33] focus on cases where the predicted parameters appear only in the objective function. In such settings, the decision loss is typically the regret, which measures suboptimality of the predicted solution. Instead, we focus on predicting parameters that appear in the constraints of the optimization problem. In this setting too, suboptimal solutions may arise, because the feasible space defined by the predicted parameters may exclude the true optimal solution. However, focusing only on suboptimality is insufficient. This is because the predicted solution may become infeasible with respect to the constraints instantiated by the true parameters. Therefore, the ML model should also be trained to minimize the likelihood of such infeasible outcomes.
@@
Focusing only on suboptimality or on infeasibility generally does not lead to a desirable balance between these two objectives. To address this, Hu et al. [14] model decision-making with uncertain constraint parameters as a two-stage process: first, a solution is generated using the predicted parameters; then, if needed, corrective actions – incurring additional costs – are carried out to resolve potential infeasibility. Subsequent works [15, 16] have extended this line of research.
@@
One limitation of their approach is that the corrective action is often problem-specific, and restricted to a specific form of optimization problem (e.g., a linear program). Instead, we avoid the complexity of a second stage altogether, and focus on predicting parameters that minimize both the infeasibility and the suboptimality of the first-stage solution itself. To this end, we first propose two novel loss functions: the Infeasibility Penalty Loss (IPL) and Optimality-Preserving Loss (OPL). The idea behind IPL is that when solving with the predicted parameters leads to a solution that is infeasible according to the true parameters, such infeasibility should be (proportionally) penalized. On the other hand, the OPL loss will penalize predicted parameters that make the true optimal solution infeasible, to avoid obtaining a suboptimal solution when solving with the predicted parameters.
@@
These two losses often contradict each other. The IPL term encourages conservative parameter predictions to guarantee feasibility, even at the expense of producing extremely suboptimal solutions. In contrast, OPL incentivizes loose parameter predictions to ensure the optimal solution is not cut away. Moreover, in practice, a decision-maker may favor more feasibility over optimality or vice versa. To support this subjective preference of one loss over the other, we propose to minimize a weighted average of the two losses. In this way, we introduce a tunable weighting mechanism that adjusts the total loss based on the decision-maker’s preference. We refer to our approach as Odece, for “optimizing decisions through end-to-end constraint estimation”.
@@
We compare Odece with the PFL approach of minimizing the mean squared error (MSE), and with some existing DFL approaches. These approaches do not allow the decision-maker to adjust the predicted solution based on their subjective preference between optimality and feasibility. Moreover, the DFL approaches for constraint prediction are limited to linear programs (LPs) or integer or mixed-integer LPs (ILPs and MILPs), or they focus on predicting full constraint systems rather than individual parameters within the constraints. Our experimental results across various optimization problem instances show that, for a single value of the tunable parameter, Odece matches or outperforms existing baselines in terms of both suboptimality and feasibility. More importantly, they show that our approach offers a simple way for decision-makers to control the trade-off between suboptimality and infeasibility using a single parameter.
@@ 2 Related Literature
Differentiating through optimization problems.
@@
Differentiating through optimization problems has gained attention in ML, as it allows solvers to be embedded into gradient-based training. This is often done by ADMM [9, 29] or by implicit differentiation [8] of the optimality conditions. Implicit differentiation techniques have been developed for quadratic programs [2] and conic optimization [1]. However, for combinatorial optimization problems like ILPs, the true gradient is zero almost everywhere because small changes in the parameters usually do not affect the optimal solution. Hence, implicit differentiation techniques, which are applicable to smooth optimization problems, cannot be applied to such combinatorial problems. To address this, Paulus et al. [26] propose CombOptNet, a technique for ILPs to compute pseudo-gradients based on the Euclidean distance between the solution and the constraints.
Predicting objective parameters.
Most DFL works focus on predicting parameters in the objective function. Some of these works are based on differentiating through the problem after smoothing it, either using regularization [13, 18, 33] or perturbation [5, 10, 25, 28]. Another approach is to define informative surrogates of the decision loss [4, 12, 19, 20, 23, 31]. Regularization-based methods require differentiable solvers, such as Cvxpylayers [1] or OptNet [2]. In contrast, perturbation and surrogate-based approaches only need access to solutions. These solutions generally do not need to come from an exact solver and can be obtained by heuristics or alternative methods. For instance, surrogate-based techniques have been applied to planning problems solved via approximate automated planners [21]. We refer readers to a recent survey paper [22] for a more comprehensive overview.
Predicting constraint parameters.
@@
Relatively few DFL papers have considered the prediction of parameters in the constraints. Most of these [14, 15, 16] model decision-making as a two-stage process. In the first stage, a solution is obtained using the predicted parameters, which may be infeasible with respect to the true parameters. The second stage then applies a corrective action to convert the first-stage solution into a feasible one. Among these works, Hu et al. [15] also recommend a strategy for applying their approach to settings where no corrective action is available at inference time. Their proposal is to assign a very high penalty factor to the corrective action during training. In contrast to this line of work, we do not introduce corrective actions even during training and focus on balancing decision quality and feasibility in a single stage. Another approach [11] simultaneously learns the objective and the constraints, but the constraints are soft and the problem must be formulated as a discrete graphical model. The work by Nandwani et al. [24] aims to learn the full set of constraints from the features. They do this by learning a set of hyperplanes that separate the optimal solution from other negative assignments. A key component of their technique is the generation of such negative assignments, which must include both infeasible as well as feasible but suboptimal ones. In contrast to their setting, we focus on learning only the parameters within a known constraint structure. Their method does not easily generalize to our setting, in which the learned parameters must remain compatible with the known constraints to make high-quality solutions. Moreover, both Hu et al. [15] and Nandwani et al. [24] restrict their methods to LPs and ILPs, while the framework proposed in this paper is more general and applicable to a broader class of optimization problems. The prediction of constraints is also central in constraint acquisition, for optimization [3, 17] or constraint satisfaction [6, 32] problems, but the setting differs because the structure of the constraints is unknown and the uncertainty does not depend on any input features.
@@ 3 Preliminaries
We will consider a constrained optimization problem (COP) of the following form:
@@
where $\Omega$ defines the domain of the decision variable vector $\bm{x}$. $\Omega$ may be a subset of real numbers or integers, depending on the type of optimization problem. The upper and lower bounds for each decision variable are also specified by $\Omega$. We will assume that the constraint functions $g_{i}(\bm{x};\bm{\rho})$ are differentiable functions. We do not explicitly express equality constraints, since any constraint of the form $h_{j}(\bm{x};\bm{\rho})=0$ can be represented by two inequality constraints: $h_{j}(\bm{x};\bm{\rho})\leq 0$ and $h_{j}(\bm{x};\bm{\rho})\leq 0$. To make the notation simpler, we denote the set of constraints by $\mathcal{S}$ and the feasible set when the constraint parameters take the value $\rho$ by $\mathcal{F}(\rho)$. More formally,
@@
We use $\bm{x}^{\star}(\bm{q},\bm{\rho})$ to denote a (possibly non-unique) optimal solution to the optimization problem for a given objective parameter $\bm{q}$ and constraint parameter $\bm{\rho}$. When multiple optimal solutions exist, we define the set of all optimal solutions as $\bm{W}(\bm{q},\bm{\rho})=\{\bm{x}:f(\bm{x};\bm{q})\leq f(\bm{x}^{\prime};\bm{q});\;\forall\bm{x}^{\prime}\in\mathcal{F}(\bm{\rho})\}$. Note that $\bm{W}(\bm{q},\bm{\rho})$ is a subset of the feasible region $\mathcal{F}(\bm{\rho})$. So, if $\mathcal{F}$ is empty (i.e., no feasible solution exists), $\bm{W}$ is also empty. In the special case of a unique optimal solution, $\bm{W}(\bm{q},\bm{\rho})$ is a singleton. However, our implementation relies on a solver that returns a single solution. When multiple optimal solutions exist, the commercial solvers we use break ties using a predefined rule.
@@
Consider the multi-dimensional 0-1 knapsack problem (MDKP) [27]. The objective of this problem is to choose a subset with maximal value from a given set of items, subject to $M$ capacity constraints. Let $\rho_{i}$ be the capacity in dimension $i$. There are $\mathit{N}$ items, the value of each item is $q_{n}$, and $\rho_{(n,i)}$ is the weight of item $n$ in dimension $i$. Each decision variable $x_{n}$ can be either zero or one, hence $\Omega=\{0,1\}^{N}$. This optimization problem can be modeled as an ILP as follows:
$\min_{x_{1:N}}\sum_{n=1}^{N}(-q_{n})x_{n}\quad\text{such that}\quad x_{n}\in\{0,1\}\;\;\forall n\in[N],\quad\sum_{n=1}^{N}\rho_{ni}x_{n}\leq\rho_{i}\;\;\forall i\in[M]$ (3)
Here, $\mathcal{S}$ consists of the $M$ capacity constraints.
@@
We will develop a framework based on maximum likelihood estimation (MLE) to predict constraint parameters. Our goal is to maximize the log-likelihood of the observed data $\mathcal{D}$. In PFL, the parameters $\theta$ of the predictive model would be estimated to increase the likelihood of the observations $\bm{\rho}_{\kappa}$ given the features $\bm{\phi}_{\kappa}$. However, in DFL training, the focus is on the likelihood of $\bm{x}^{\star}(\bm{q}_{\kappa},\bm{\rho}_{\kappa})$ being the optimal solution, given $\bm{q}_{\kappa}$ and $\bm{\phi}_{\kappa}$. To put it another way, the MLE perspective of DFL consists of learning a $\theta$ to predict the constraint parameter vector $\bm{\hat{\rho}}_{\kappa}$ so that $\bm{x}^{\star}(\bm{q}_{\kappa},\bm{\rho}_{\kappa})$ is the most likely solution to the parameter $\bm{\hat{\rho}}_{\kappa}=\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})$.
@@ 4 Methodology
We start by defining a function $\mathit{UNSAT}_{i}(\bm{x},\bm{\rho}_{\kappa})$ to indicate whether the $i$-th constraint is violated by an assignment $\bm{x}$ when the constraint parameter is set to the vector $\bm{\rho}_{\kappa}$. It is defined as:
@@
As we want to maximize the likelihood of obtaining a feasible solution with optimal objective value, we first model the probability of violating the constraints using the predicted parameters. Given predicted parameters $\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})$, we model the parametric probability $\mathit{P}_{\theta}\big(\mathit{UNSAT}_{i}(\bm{x},\mathcal{M}_{\theta}(\bm{\phi}_{\kappa}))\big|\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\big)$ in the following manner:
@@
As $\mathit{P}_{\theta}\left(\mathit{SAT}_{i}\left(\bm{x},\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\right)\middle|\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\right)$ = 11- $\mathit{P}_{\theta}\left(\mathit{UNSAT}_{i}\left(\bm{x},\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\right)\middle|\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\right)$, we can write:
@@
4.1 Loss Function to Penalize Infeasibility-Inducing Predictions
We first propose a loss function that ensures that true infeasible assignments remain infeasible under the predicted parameters. Let $\bm{x}_{\mathit{neg}}$ be an assignment that is infeasible under the true parameters. As it is easy to determine which constraints are violated by $\bm{x}_{\mathit{neg}}$, we develop an MLE framework to maximize the likelihood of violating those constraints under the predicted parameters. Given a $\bm{x}_{\mathit{neg}}$, the log-likelihood of violating these constraints can be written in the following form: